# Action2Sound: Ambient-Aware Generation of Action Sounds from Egocentric Videos

### v2

@@ Abstract
Generating realistic audio for human interactions is important for many applications, such as creating sound effects for films or virtual reality games. Existing approaches implicitly assume total correspondence between the video and audio during training, yet many sounds happen off-screen and have weak to no correspondence with the visuals—resulting in uncontrolled ambient sounds or hallucinations at test time. We propose a novel ambient-aware audio generation model, AV-LDM. We devise a novel audio-conditioning mechanism to learn to disentangle foreground action sounds from the ambient background sounds in in-the-wild training videos. Given a novel silent video, our model uses retrieval-augmented generation to create audio that matches the visual content both semantically and temporally. We train and evaluate our model on two in-the-wild egocentric video datasets Ego4D and EPIC-KITCHENS. Our model outperforms an array of existing methods, allows controllable generation of the ambient sound, and even shows promise for generalizing to computer graphics game clips. Overall, our work is the first to focus video-to-audio generation faithfully on the observed visual content despite training from uncurated clips with natural background sounds.
@@
audio-visual learning egocentric video understanding action sound generation
@@ 1 Introduction
We interact with objects around us in our daily lives and these actions often produce sound as a result of physical interactions, e.g., clicking on a mouse, closing a door, or cutting vegetables. The distinct characteristics of these action sounds depend upon the type of action being performed, the shape and material composition of the objects being acted upon, the amount of force being applied, and so forth. Vision not only captures what physical interaction happens but also informs us when the interaction happens, suggesting the possiblity of synthesizing semantically plausible and temporally synchronous action sounds from silent videos alone. This capability would accelerate many real-world applications, such as text-to-video generation, generating sound effects for films (Foley), or sound effect generation for virtual reality (VR) and video games.
@@
Figure 1: Real-world audio consists of both foreground action sounds (whose causes are visible in the FoV) and background ambient sounds that are generated by sources offscreen. Whereas prior work is agnostic to this division when performing generation, our method is ambient-aware and disentangles action sound from ambient sound. Our key technical insight is how to train with in-the-wild videos exhibiting natural ambient sounds, while still learning to factor out their effects on generation. The green arrows reference how we condition generation on sound from a related, but time-distinct, video clip to achieve this.
@@
Some prior work studies impact sound synthesis from videos [42, 48] while others target more general video-to-audio generation [25, 38]. All these methods implicitly assume total correspondence between the video and audio and aim to generate the whole target audio from the video. However, this strategy falls short for in-the-wild training videos, which are rife with off-screen ambient sounds, e.g., traffic noise, people talking, or A/C running. While some of these ambient sounds are weakly correlated with the visual scene, such as the wind blowing in an outdoor environment, many of them have no visual correspondence, such as off-screen speech or a stationary buzzing noise from the fridge. Existing methods are not able to disentangle action sounds from ambient sounds and treat them as a whole, leading to uncontrolled generation of ambient sounds at test time and sometimes even hallucination, e.g., random action or ambient sounds. This is particularly problematic for generating action sounds because they are often subtle and transient compared to the ambient sounds. For example, trained in the traditional way, a model given a scene that looks like a noisy restaurant risks generating “restaurant-like" ambient sounds, while ignoring the actual movements and activities of the foreground actions, such as a person stirring their coffee with a metal spoon.
@@
How can we disentangle the foreground action sounds from background ambient sounds for in-the-wild video data without ground truth separated streams? Simply applying a noise removal algorithm on the target audio does not work well since in-the-wild blind source separation of general sounds from a single microphone is still an open challenge [49]. The key observation we have is that while action sounds are highly localized in time, ambient sounds tend to persist across time. Given this observation, we propose a simple but effective solution to disentangle ambient and action sounds: during training, in addition to the input video clip, we also condition the generation model on an audio clip from the same long video as the input video clip but from different timestamps. See Fig. 1. By doing so, we lift the burden of generating energy-dominating ambient sounds and encourage the model to focus on learning action cues from the visual frames to generate action sounds. At test time, we do not assume access to (even other clips of) the ground truth video/audio. Instead, we propose to retrieve an audio segment from the training set with an audio-visual similarity scoring model, inspired by recent ideas in retrieval-augmented generation (RAG) [29, 19, 34]. This benefits examples where the visual scene has a weak correlation with the ambient sound that is appealing to capture, e.g., outdoor environments.
@@
Existing action sound generation work relies on either clean, manually-collected data that has a limited number of action categories [42, 48, 8], or videos crawled from YouTube based on predefined taxonomies [15, 5, 25]. To expand the boundary of action sound generation to in-the-wild human actions, we take advantage of recent large-scale egocentric video datasets [18, 9]. Though our model is not tailored to egocentric video in any way, there are two main benefits of using these datasets: 1) egocentric videos provide a close view of human actions compared to exocentric videos, where hand-object interactions are much smaller from a distance and often occluded, and 2) these datasets have timestamped narrations describing atomic actions. We design an automatic pipeline to extract and process clips from Ego4D, and curate Ego4D-Sounds with 1.2 million audio-visual action clips.
@@
Our idea of disentangling action and ambient sounds implicitly in training is model-agnostic. In this paper, we instantiate it by designing an audio-visual latent diffusion model (AV-LDM) that conditions on both modality streams for audio generation. We evaluate our AV-LDM against recent works on a wide variety of metrics and show that our model outperforms the existing methods significantly on both Ego4D-Sounds and EPIC-KITCHENS. We conduct a human evaluation study that shows our model synthesizes plausible action sounds according to the video. Please see/listen for yourself in our supplementary video! We also show promising preliminary results on virtual reality game clips. To the best of our knowledge, this is the first work that demonstrates the disentanglement of foreground action sounds from background sounds for action-to-sound generation on in-the-wild videos.
@@ 2.1 Action Sound Generation
A pioneering work for capturing human-generated action sounds collects videos where people hit, scratch, or prod objects with a drumstick [42]. This is an early inspirational effort, though it is by design limited in the type of actions. The robotics community also studies this problem by using robotic platforms to collect collision sounds and analyze or synthesize them from video [14, 8]. Other works approach this problem by building a simulation for collision events [13]; however, it is hard for computational approaches to simulate the impact or general action sounds due to the complexity of the physical interactions. Existing methods demonstrate good synthesis results when the data are noise-free. However, they are not equipped to learn from in-the-wild action videos, where the action sound is always coupled with ambient sound. We propose an ambient-aware model to deal with this issue head-on and also introduce the Ego4D-Sounds dataset to expand action sound synthesis to in-the-wild actions.
@@ 2.2 Egocentric Video Understanding with Audio
Understanding human activities in videos has long been a core challenge of computer vision. Early research studies activity recognition from exocentric videos such as UCF101 [47], Kinetics [27], or ActivityNet [12]. Recent work explores the egocentric setting and introduces large egocentric datasets such as Ego4D [18] or EPIC-KITCHENS [9]. Leveraging both the video and audio streams in egocentric videos, many interesting tasks are enhanced, such as action recognition [28], localization [44], active speaker localization [26], sounding object localization [22], and state-aware visual representations from audible interactions [39]. Most related to our work is SoundingActions [4] that learns visual representations of actions that make sounds, which is valuable for indexing and recognition problem settings, but ill-equipped for generation, as we show later when retrieval models [16] perform poorly for generating synchronized audio. All existing audio-visual learning for egocentric video focuses on perception, i.e., understanding what happens in the video. In contrast, we target the video-to-audio generation problem. Furthermore, relative to any of the above, our idea to implicitly learn to disentangle the action sound from ambient sounds is novel.
@@ 2.3 Diffusion Models and Conditional Audio Generation
Diffusion models have attracted significant attention recently because of their high fidelity generation [10, 40, 45]. Initially proposed for image generation [46, 20], dthey have also been successfully applied to speech and audio generation [33, 43, 51, 36, 23]. Benefitting from classifier-free guidance [21] and large-scale representation learning, AudioLDM [36] and Make-An-Audio [23] have demonstrated successful diffusion-based text-to-audio generation. More recently, Diff-Foley [38] adapts latent diffusion models for video-to-audio generation by first conducting audio-video contrastive learning and then video-conditioned audio generation. While this approach demonstrates promising results, it does not address the background ambient sound problem. Inspired by recent work on retrieval-augmented generation (RAG) for text  [29, 19, 34, 3] and image generation [6, 2], we show how our audio-conditioning insight carries over to inference time via a retrieval component of the model. Conditional video-to-audio generation conditions on either a physics prior to guide diffusion-based impact sound generation [48] or, in CondFoleyGen [11], another video clip to modify the characteristics of the action sound. Our method also considers additional conditioning signals to control the output, but for a very different purpose; our model is the first to address foreground/background sound disentanglement in generation.
@@ 3 Ambient-aware Action Sound Generation
We first discuss our high-level idea of how to guide the generation model to disentangle action sounds from ambient sounds. We then extend the latent diffusion models (LDM) to accommodate both audio and video conditions, which we name AV-LDM. We also discuss our pretraining stage.
@@ 3.1 Action-to-Sound Generation
Given a video $V\in\mathbb{R}^{(T*S_{V})\times H\times W\times 3}$, where $T$ is the duration of the video and $S_{V}$ is the video sample rate, and the accompanying audio waveform $A\in\mathbb{R}^{1\times(T*S_{A})}$, where $S_{A}$ is the audio sample rate, our goal is to model the conditional distribution $p(A|V)$ for video-to-audio generation. During training we observe natural video coupled with its audio, whereas at inference time we have only a silent video—e.g., could be an output from text-to-video generation, or a VR/video game clip, or simply a real-world video for which we want to generate new plausible sounds.
@@ 3.2 Disentangling Action and Ambient Sounds
Therefore, it’s beneficial to proactively disentangle action sounds and ambient sounds during training. However, separating in-the-wild ambient sounds is still an open challenge as recent models rely on supervised training on artificially mixed sounds, for which the ground truth complex masks can be obtained [49]. Simply applying off-the-shelf noise reduction methods to training data leads to poor performance, as we will show in Sec. 5.
While it is difficult to explicitly separate the ambient and action sound in the target audio, our key observation is that ambient sounds are usually fairly stationary across time. Given this observation, we propose a simple but effective method to achieve the disentanglement. During training, in addition to video clip $V$, we also provide the model an audio clip $A_{n}$ that comes from the same training video but a different timestamp as the input video clip (see Fig. 3). Therefore, instead of modeling $p(A|V)$, we model $p(A|V,A_{n})$. Given the hypothesis that $A_{n}$ is likely to share ambient sound characteristics with $A$, it can take away the burden of learning weakly correlated or even uncorrelated ambient sounds from visual input alone, and encourages the model to focus on learning action features from the visual input. For the selection of $A_{n}$, we randomly sample one audio clip from the nearest $X$ clips in time. While there is no guarantee that the sampled audio shares exactly the same ambient sound with the target audio, their ambient sounds should largely overlap since they are close in time, which provides a consistent learning signal to help the model learn the disentanglement.
@@ 3.3 Retrieval Augmented Generation and Controllable Generation
While during training we have access to the clips in the same long video as the input clip, we of course cannot access that information at test time. How we select $A_{n}$ at test time depends on the purpose of the generation. We consider two use cases: action-ambient joint generation and action-focused generation. In the first scenario, we would like the model to generate both the action sound and the ambient sound that is plausible for the visual environment. This is, for example, useful for generating sound effects for videos. In the latter scenario, we would like the model to focus the generation on action sounds and minimize ambient sounds, which is useful, for example, for generating sounds for games. Fig. 4 depicts the two scenarios.
@@
For action-focused generation, we want $A_{n}$ to have minimal ambient level. We find simply filling $A_{n}$ with all zeros results in poor performance, likely because it is too far out of the training distribution. Instead, we find conditioning the generation on a low-ambient sound will hint the model to focus on action sound generation and generate minimal ambient sound. See Sec. 5.3.
Figure 4: Two inference settings: “action-ambient joint generation” and “action-focused generation”. In the first setting, we condition on audio retrieved from the training set and aim to generate both plausible action and ambient sounds. In the second setting, we specify an audio file with low ambient sound and the model focuses on generating plausible action sounds while minimizing the ambient sounds.
@@ 4 The Ego4D-Sounds Dataset
Figure 5: Example clips in Ego4D-Sounds. We show one video frame, the action description, and the sound for each example. Note how these actions are subtle and long-tail, usually not present in typical
video datasets.
@@
Next we describe our efforts to curate Ego4D-Sounds, an audio-video dataset for human action sound generation. Our goal is to curate a high-quality dataset for action-audio correspondence for action-to-sound generation, addressing the issue of limited action types in the existing impact sound datasets [42, 7].
Ego4D [18] is an existing large-scale egocentric video dataset that has more than 3,600 hours of video recordings depicting hundreds of daily activities; 2,113 of those hours have audio available. It also has time-stamped narrations that are free-form sentences describing the current activity performed by the camera-wearer. We first utilize the narration timestamps in Ego4D to extract clips. However, not all clips have meaningful action sounds and there are many actions like “talk with someone", “look around", “turn around" that have low audio-visual correspondence. We then use an automatic pipeline to process all extracted clips to create the Ego4D-Sounds dataset, which has 1.2 million audio-visual action clips. Similarly, for the test set, we curate 11k clips for evaluation. See Supp. for more details on the data processing pipeline. We show examples in Fig. 5 and comparison with other datasets in Tab. 1.

### v3

@@ Abstract
Generating realistic audio for human actions is important for many applications, such as creating sound effects for films or virtual reality games. Existing approaches implicitly assume total correspondence between the video and audio during training, yet many sounds happen off-screen and have weak to no correspondence with the visuals—resulting in uncontrolled ambient sounds or hallucinations at test time. We propose a novel ambient-aware audio generation model, AV-LDM. We devise a novel audio-conditioning mechanism to learn to disentangle foreground action sounds from the ambient background sounds in in-the-wild training videos. Given a novel silent video, our model uses retrieval-augmented generation to create audio that matches the visual content both semantically and temporally. We train and evaluate our model on two in-the-wild egocentric video datasets, Ego4D and EPIC-KITCHENS, and we introduce Ego4D-Sounds—1.2M curated clips with action-audio correspondence. Our model outperforms an array of existing methods, allows controllable generation of the ambient sound, and even shows promise for generalizing to computer graphics game clips. Overall, our approach is the first to focus video-to-audio generation faithfully on the observed visual content despite training from uncurated clips with natural background sounds.
@@
audio-visual learning egocentric video understanding
@@ 1 Introduction
As we interact with objects around us in our daily lives, our physical actions often produce sound, e.g., clicking on a mouse, closing a door, or cutting vegetables. The distinct characteristics of these action sounds depend upon the type of action being performed, the shapes and materials of the objects being acted upon, the amount of force being applied, and so forth. Vision not only captures what physical interaction happens but also informs us when the interaction happens, suggesting the possiblity of synthesizing semantically plausible and temporally synchronous action sounds from silent videos alone. This capability would accelerate many real-world applications, such as text-to-video generation, generating sound effects for films (Foley), or sound effect generation for virtual reality (VR) and video games.
@@
Figure 1: Real-world audio consists of both foreground action sounds (whose causes are visible) and background ambient sounds generated by sources offscreen. Whereas prior generation work is agnostic to this division, our method is ambient-aware and disentangles action sound from ambient sound. Our key technical insight is how to train with in-the-wild videos exhibiting natural ambient sounds, while still learning to factor out their effects on generation. The green arrows reference how we condition generation on sound from a related, but time-distinct, video clip to achieve this.
@@
Some prior work studies impact sound synthesis from videos [46, 52] while others target more general video-to-audio generation [28, 41]. These methods implicitly assume total correspondence between the video and audio and aim to generate the whole target audio from the video. However, this strategy falls short for in-the-wild training videos, which are rife with off-screen ambient sounds, e.g., traffic noise, people talking, or A/C running. While some of these ambient sounds are weakly correlated with the visual scene, such as the wind blowing in an outdoor environment, many of them have no visual correspondence, such as off-screen speech or a stationary buzzing noise from the fridge. Most existing methods are not able to disentangle action sounds from ambient sounds and treat them as a whole, leading to uncontrolled generation of ambient sounds at test time and sometimes even hallucination, e.g., random action or ambient sounds. This is particularly problematic for generating action sounds because they are often subtle and transient compared to ambient sounds. For example, trained in the traditional way, a model given a scene that looks like a noisy restaurant risks generating “restaurant-like” ambient sounds, while ignoring the actual movements and activities of the foreground actions, such as a person stirring their coffee with a metal spoon.
How can we disentangle the foreground action sounds from background ambient sounds for in-the-wild video data without ground truth separated streams? Simply applying a noise removal algorithm on the target audio does not work well since in-the-wild blind source separation of general sounds from a single microphone is still an open challenge [53], and class-dependent models for predicting visually relevant sounds cannot generalize to in-the-wild video [7].
@@
Our key observation is that while action sounds are highly localized in time, ambient sounds tend to persist across time. Given this observation, we propose a simple but effective solution to disentangle ambient and action sounds: during training, in addition to the input video clip, we also condition the generation model on an audio clip from the same long video as the input video clip but from different timestamps. See Fig. 1. By doing so, we lift the burden of generating energy-dominating ambient sounds and encourage the model to focus on learning action cues from the visual frames to generate action sounds. At test time, we do not assume access to (even other clips of) the ground truth video/audio. Instead, we propose to retrieve an audio segment from the training set with an audio-visual similarity scoring model, inspired by recent ideas in retrieval-augmented generation (RAG) [32, 22, 37]. This benefits examples where the visual scene has a weak correlation with the ambient sound that is beneficial to capture, e.g., outdoor environments.
@@
Existing action sound generation work relies on either clean, manually-collected data that has a limited number of action categories [46, 52, 10], or videos crawled from YouTube based on predefined taxonomies [17, 6, 28, 7]. To expand the boundary of action sound generation to in-the-wild human actions, we take advantage of recent large-scale egocentric video datasets [20, 11]. Though our model is not tailored to egocentric video in any way, there are two main benefits of using these datasets: 1) egocentric videos provide a close view of human actions compared to exocentric videos, where hand-object interactions are observed from a distance and are often occluded, and 2) these datasets have timestamped narrations describing atomic actions. We design a pipeline to extract and process clips from Ego4D, curating Ego4D-Sounds with 1.2 million audio-visual action clips. [^1] https://ego4dsounds.github.io
@@
Our idea of disentangling action and ambient sounds implicitly in training is model-agnostic. In this paper, we instantiate it by designing an audio-visual latent diffusion model (AV-LDM) that conditions on both modality streams for audio generation. We evaluate our AV-LDM against recent work on a wide variety of metrics and show that our model outperforms the existing methods significantly on both Ego4D-Sounds and EPIC-KITCHENS. We conduct a human evaluation study that shows our model synthesizes plausible action sounds according to the video. Please see/listen for yourself in our supplementary video! We also show promising preliminary results on virtual reality game clips. To the best of our knowledge, this is the first work that demonstrates the disentanglement of foreground action sounds from background sounds for action-to-sound generation on in-the-wild videos.
@@ 2.1 Action Sound Generation
A pioneering work for capturing human-generated action sounds collects videos where people hit, scratch, or prod objects with a drumstick [46]. This is an early inspirational effort, though it is by design limited in the type of actions. The robotics community also studies this problem by using robotic platforms to collect collision sounds and analyze or synthesize them from video [16, 10]. Other work simulates collision events [15], which remains difficult for action sounds due to the complexity of the physical interactions. Most existing methods demonstrate good synthesis results when the data are noise-free. However, they are not equipped to learn from in-the-wild action videos, where the action sound is always coupled with ambient sound. Sharing our motivation to disregard irrelevant sounds, the REGNET framework [7] aims to predict visually relevant sounds by conditioning on ground truth audio with a bottleneck design. However, it does not allow controllable generation and risks learning to copy the target action and ambient sound leading to weaker empirical performance. More importantly, REGNET [7] requires curated datasets to train class-dependent models, which prevents generalization to in-the-wild data, as we will see in results. We propose an ambient-aware model to deal with this issue head-on and also introduce the Ego4D-Sounds dataset to expand action sound synthesis to in-the-wild actions.
@@ 2.2 Egocentric Video Understanding with Audio
Understanding human activities in videos has long been a core challenge of computer vision. Early research studies activity recognition from exocentric video [51, 30, 14]. Recent work explores the egocentric setting and introduces large egocentric datasets such as Ego4D [20], EPIC-KITCHENS [11], and Ego-Exo4D [21]. Leveraging both the video and audio streams in egocentric videos, many interesting tasks are enhanced, such as action recognition [31], localization [48], active speaker localization [29, 42], sounding object localization [25, 5], and state-aware representations [43]. Most related to our work is SoundingActions [4] which learns visual representations of actions that make sounds, and is valuable for indexing and recognition problem settings, but ill-equipped for generation, as we show later. All existing audio-visual learning for egocentric video focuses on perception, i.e., understanding what happens in the video. In contrast, we target the video-to-audio generation problem. Furthermore, relative to any of the above, our idea to implicitly learn to disentangle the action sound from ambient sounds is novel.
@@ 2.3 Diffusion Models and Conditional Audio Generation
Diffusion models have attracted significant attention recently because of their high fidelity generation [12, 44, 49]. Initially proposed for image generation [50, 23], they have also been successfully applied to speech and audio generation [36, 47, 55, 39, 26]. Benefitting from classifier-free guidance [24] and large-scale representation learning, AudioLDM [39] and Make-An-Audio [26] perform diffusion-based text-to-audio generation. More recently, Diff-Foley [41] adapts latent diffusion models for video-to-audio generation by first conducting audio-video contrastive learning and then video-conditioned audio generation. While promising, it does not address the background ambient sound problem. Inspired by recent work on retrieval-augmented generation (RAG) for text [32, 22, 37, 3] and image generation [8, 2], we show our audio-conditioning insight carries over to inference time via a retrieval component of the model. Conditional video-to-audio generation conditions on either a physics prior to guide diffusion-based impact sound generation [52] or, in CondFoleyGen [13], another video clip to modify characteristics of the action sound. Our method also considers additional conditioning signals to control the output, but for a very different purpose; our model is the first to address foreground/background sound disentanglement in generation.
@@ 3 Ambient-aware Action Sound Generation
We first discuss our high-level idea of how to guide the generation model to disentangle action sounds from ambient sounds. We then devise AV-LDM, an extension of latent diffusion models (LDM) to accommodate both audio and video conditions. We also discuss our pretraining stage.
@@ 3.1 Action-to-Sound Generation
Given a video $V\in\mathbb{R}^{(T*S_{V})\times H\times W\times 3}$, where $T$ is the duration of the video and $S_{V}$ is the video sample rate, and the accompanying audio waveform $A\in\mathbb{R}^{1\times(T*S_{A})}$, where $S_{A}$ is the audio sample rate, our goal is to model the conditional distribution $p(A|V)$ for video-to-audio generation. During training we observe natural video coupled with its audio, whereas at inference time we have only a silent video—e.g., could be an output from text-to-video, or a VR/video game clip, or simply a real-world video for which we want to generate new plausible sounds.
@@ 3.2 Disentangling Action and Ambient Sounds
Therefore, it is important to proactively disentangle action sounds and ambient sounds during training. However, separating in-the-wild ambient sounds is still an open challenge: recent models rely on supervised training using artificially mixed sounds, for which the ground truth complex masks can be obtained [53]. Simply applying off-the-shelf noise reduction methods to training data leads to poor performance, as we will show in Sec. 5.
While it is difficult to explicitly separate the ambient and action sound in the target audio, our key observation is that ambient sounds are usually fairly stationary across time. Given this observation, we propose a simple but effective method to achieve the disentanglement. During training, in addition to video clip $V$, we also provide the model an audio clip $A_{n}$ that comes from the same training video but a different timestamp as the input video clip (see Fig. 3). Therefore, instead of modeling $p(A|V)$, we model $p(A|V,A_{n})$. Given the hypothesis that $A_{n}$ is likely to share ambient sound characteristics with $A$, it can take away the burden of learning weakly correlated or even uncorrelated ambient sounds from visual input alone, and encourages the model to focus on learning action features from the visual input. For the selection of $A_{n}$, we randomly sample one audio clip from the nearest $X$ clips in time. While there is no guarantee that the sampled audio shares exactly the same ambient sound with the target audio, their ambient sounds should largely overlap since they are close in time, which provides a consistent learning signal to help the model learn the disentanglement. While is possible for the sampled audio to contain repetitions of the target action sound, 1) the chance of selecting a semantically relevant sound low (9% based on (verb,noun) taxonomy) and 2) the precise temporal onset is almost never the same, thus making it impossible for the model to cheat in training.
@@ 3.3 Retrieval Augmented Generation and Controllable Generation
While during training we have access to the clips in the same long video as the input clip, we of course cannot access that information at test time. How we select $A_{n}$ at test time depends on the purpose of the generation. We consider two use cases: action-ambient joint generation and action-focused generation. In the first scenario, we would like the model to generate both the action sound and the ambient sound that is plausible for the visual environment. This is useful, for example, for generating sound effects for videos. In the latter scenario, we would like the model to focus the generation on action sounds and minimize ambient sounds, which is useful, for example, for generating sounds for games. Fig. 4 depicts the two scenarios.
@@
For action-focused generation, we want $A_{n}$ to have minimal ambient level. We find simply filling $A_{n}$ with all zeros results in poor performance, likely because it is too far out of the training distribution. Instead, we find conditioning the generation on a low-ambient sound will cue the model to focus on action sound generation and generate minimal ambient sound. See Sec. 5.2.
Figure 4: Two inference settings: “action-ambient joint generation” and “action-focused generation”. In the first setting, we condition on audio retrieved from the training set and aim to generate both plausible action and ambient sounds. In the second setting, we condition on an audio file with low ambient sound and the model focuses on generating plausible action sounds while minimizing the ambient sounds.
@@ 4 The Ego4D-Sounds Dataset
Figure 5: Example clips in Ego4D-Sounds. We show one video frame, the action description, and the sound for each example. Note how these actions are subtle and long-tail, usually not present in typical video datasets.
@@
Next we describe how we curate Ego4D-Sounds, an audio-video dataset for human action sound generation. Our goal is to curate a large-scale high-quality dataset for action-audio correspondence for action-to-sound generation, addressing the issue of limited action types and scale in the existing impact sound datasets [46, 9], as well as more general audio-video datasets [6, 27].
Ego4D [20] is an existing large-scale egocentric video dataset that has more than 3,600 hours of video recordings depicting hundreds of daily activities; 2,113 of those hours have audio available. It also has time-stamped narrations that are free-form sentences describing the current activity performed by the camera-wearer. Utilizing the narration timestamps in Ego4D to extract clips directly results in a noisy dataset, since not all clips have meaningful action sounds and there are many actions like “talk with someone”, “look around”, “turn around” that have low audio-visual correspondence. To ensure Ego4D-Sounds has high action-sound correspondence, we use an automatic pipeline that consists of metadata-based filtering, audio tagging, and energy-based filtering to process all extracted clips, which yields 1.2 million audio-visual action clips plus 11k clips for evaluation. See Supp. for more details on the data processing pipeline. We show examples in Fig. 5 and comparison with other datasets in Tab. 1.